OpenStudy (anonymous):

If the slope of a line is 5/3 and it passes through points (6,4) and (x,8) what is the value of x? I'm at y-4=5/3(x-6) but I'm sorta stuck. Someone please help!:)

OpenStudy (amistre64):

let y=8 :/

OpenStudy (amistre64):

(y0-y1)/(x0-x1) = 5/3 (8-4)/(x-6) = 5/3 4/(x-6) = 5/3 12 = 5 (x-6) etc ...

OpenStudy (anonymous):

Thank you